the distance of the point (1,1,1) from the plane passing through the point(-1,-2,-1) and whose normal is perpendicular to both the lines L1 and L2 is?

Grade:11

1 Answers

Arun
3 years ago
The given equations of line are 
 
L1 : 3x+1 = 1
y+2 = 2
z+1 and L 2 : 1
x−2 = 2
y+2 = 3
z−3
Plane is given by (x+1)−7(y+2)+5(z+1)=0
 
⇒x+7y−5z+10=0
 
⇒ distance = 
75
1+7−5+10 = 75
 
13